# Compositional simulation-based inference for time series

### v1

@@ abstract
Amortized simulation-based inference (SBI) methods train neural networks on simulated data to perform Bayesian inference. While this approach avoids the need for tractable likelihoods, it often requires a large number of simulations and has been challenging to scale to time-series data. Scientific simulators frequently emulate real-world dynamics through thousands of single-state transitions over time. We propose an SBI framework that can exploit such Markovian simulators by locally identifying parameters consistent with individual state transitions. We then compose these local results to obtain a posterior over parameters that align with the entire time series observation. We focus on applying this approach to neural posterior score estimation but also show how it can be applied, e.g., to neural likelihood (ratio) estimation. We demonstrate that our approach is more simulation-efficient than directly estimating the global posterior on several synthetic benchmark tasks and simulators used in ecology and epidemiology. Finally, we validate scalability and simulation efficiency of our approach by applying it to a high-dimensional Kolmogorov flow simulator with around one million dimensions in the data domain.
@@ 1 Introduction
Numerical simulations are a central approach for tackling problems in a wide range of scientific and engineering disciplines, including physics (Brehmer & Cranmer, 2022; Dax et al., 2021), molecular dynamics (Hollingsworth & Dror, 2018), neuroscience (Gonçalves et al., 2020) and climate science (Watson-Parris et al., 2021). Simulators often include at least some parameters that cannot be measured experimentally. Inferring such parameters from observed data is a fundamental challenge. Bayesian inference provides a principled approach to identifying parameters that align with empirical observations. Standard algorithms for Bayesian inference, such as Markov Chain Monte Carlo (MCMC) (Gilks et al., 1995) and variational inference (Beal, 2003), generally require access to the likelihoods $p({\bm{x}}|{\bm{\theta}})$. However, for many simulators, directly evaluating the likelihood remains intractable, rendering conventional Bayesian approaches inapplicable. However, generating synthetic data ${\bm{x}}\sim p({\bm{x}}|{\bm{\theta}})$ is feasible for numerical simulators. Simulation-based inference (SBI) methods offer a powerful alternative to perform Bayesian inference for such simulator models with intractable likelihoods (Cranmer et al., 2020).
@@
Fig 1: Illustrative figure of the methodology: Factorized Neural Score Estimation (FNSE) method. Here, $a$ denotes the diffusion time, and ${\bm{\theta}}_{a}$ is the associated noisy parameter. The training process involves single-state transitions starting from arbitrary proposal $\tilde{p}({\bm{x}}^{t})$, with different parameters sampled from a prior distribution. During inference, the time series is divided into single state transitions, and each state transition is evaluated by the neural network to estimate local posterior scores. These local estimates are then aggregated to form a global approximation, which is subsequently used to sample from the overall posterior distribution.
@@
Classical SBI methods, such as Approximate Bayesian Computation (ABC) (Beaumont et al., 2002) and synthetic likelihoods (Wood, 2010), struggle to effectively scale to high-dimensional simulations. To address this, SBI methods using neural networks have been developed, which train networks to represent likelihoods (Papamakarios et al., 2019; Glöckler et al., 2022; Boelts et al., 2022), likelihood ratios (Durkan et al., 2020; Hermans et al., 2020; 2022; Miller et al., 2022), posteriors (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Deistler et al., 2022; Geffner et al., 2023; Wildberger et al., 2023; Sharrock et al., 2024) or target several properties at once (Radev et al., 2023; Gloeckler et al., 2024). These methods allow for parameter inference without requiring additional simulations after training, making them more efficient than traditional approaches (Lueckmann et al., 2021): they effectively amortize the cost of the simulation, and/or the full inference approach.
@@
However, applying these neural amortized methods to time series simulations can be challenging due to the high computational cost of repeated simulator calls. Running numerous time series simulations—possibly with varying sequence lengths—can be computationally prohibitive or at least wasteful. Instead, it seems advantageous to exploit the temporal structure of these simulators. For example, many scientific simulators are based on solving (stochastic) differential equations that model real-world time series dynamics through state transitions over time. Consequently, many time series simulators have an inherently Markovian structure, which can be leveraged for efficient inference. In particular, we can tackle the problem on the level of single-state transitions, a simpler task, that can be tackled with fewer simulations.
@@
In this work, we propose a framework for efficient simulation-based inference for Markovian time series simulators. Unlike other neural SBI methods that require long sequences for training, our approach locally estimates parameters consistent with a single state transition. By aggregating these local solutions, we recover the global posterior for time series of arbitrary length, significantly reducing simulation costs (Fig. 1). In this work, we focus on applying this framework to posterior score-based approaches and additionally apply it to likelihood(-ratio) estimation methods. A related challenge has been addressed in prior work  (Geffner et al., 2023; Linhart et al., 2024; Boelts et al., 2022) which has focused on efficient inference on i.i.d. (independently and identically distributed) observations. We empirically evaluate our approach on a series of synthetic benchmark tasks, as well as simulators from ecology and epidemiology. Finally, we demonstrate the superior performance of this approach on high-dimensional tasks, such as the Kolmogorov flow problem, with improved simulation efficiency and scalability.
@@ 2.1 Problem setting
To simplify the exposition, we constrain the initial condition to be independent of the parameter, $p({\bm{x}}^{0}|{\bm{\theta}})=p({\bm{x}}^{0})$, and assume that the simulation transition $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ does not depend on time $t$. We relax both of these assumptions in Appendix A.2, and also address the general case of higher-order Markov chains in Appendix A.2.3. We assume one can sample from the state-transition function ${\bm{x}}^{t+1}\sim\mathcal{T}({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$, but not to its probability density.
@@ 2.2 Amortized methods for simulation-based inference
Neural Posterior Estimation (NPE): Neural Posterior Estimation (NPE; (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Wildberger et al., 2023)) methods typically use a conditional neural density estimator, e.g., a normalizing flow (Rezende & Mohamed, 2015; Papamakarios et al., 2017; Durkan et al., 2019) to approximate the target posterior distribution. The model $q_{\phi}({\bm{\theta}}|{\bm{x}}^{0:T})$ is trained via maximum likelihood to estimate $p({\bm{\theta}}|{\bm{x}}^{0:T})$ from a dataset of parameter-data pairs $\{({\bm{\theta}}_{i},{\bm{x}}_{i}^{0:T})\}_{i=1}$, where $({\bm{\theta}}_{i},{\bm{x}}_{i}^{0:T})\sim p({\bm{\theta}})p({\bm{x}}^{0:T}|{\bm{\theta}})$. For observations of inhomogeneous size, NPE often incorporates an embedding network. Embedding networks can be designed as recurrent networks (Lueckmann et al., 2017), permutation-invariant architectures (Chan et al., 2018; Radev et al., 2020), or attention-based models (Schmitt et al., 2023) to respect known invariances in the data. At inference time, NPE yields an amortized approximation of the posterior for any observation ${\bm{x}}_{o}^{0:T}$ therefore offering rapid evaluations or samples of the posterior.
Neural likelihood(-ratio) estimation (NLE, NRE): Neural Likelihood(-Ratio) Estimation (NLE (Papamakarios et al., 2019), NRE (Hermans et al., 2020; 2022; Miller et al., 2022)) methods do not directly approximate the posterior distribution but instead train a surrogate $q_{\psi}({\bm{x}}_{i}^{0:T}|{\bm{\theta}})$ for the likelihood $p({\bm{x}}_{i}^{0:T}|{\bm{\theta}})$ or likelihood ratio $p({\bm{x}}_{i}^{0:T}|{\bm{\theta}})/p({\bm{x}}_{i}^{0:T})$. Once the likelihood is accessible via the surrogate, standard techniques such as MCMC or variational inference can be applied to perform inference. These approacehs circumvent the problem of not having access to the likelihood, but still leaves us with the drawbacks of standard inference methods, i.e., slow sampling and potential failure modes such as robustly handling multimodality (Glöckler et al., 2022). Nonetheless, they also inherit the flexibility of standard inference techniques, such as the ability to handle multiple independent (or conditionally independent) observations by simply multiplying the likelihood terms (Boelts et al., 2022).
@@
This recently proposed approach has been highly successful across various tasks (Geffner et al., 2023; Wildberger et al., 2023; Sharrock et al., 2024; Gloeckler et al., 2024). It offers a trade-off between the efficiency of the static NPE method and the flexibility of slower but more flexible NL(R)E method. By enabling feasible post-hoc modifications and composability through appropriate adjustments to the backward diffusion process, it bridges the gap between these two approaches (Geffner et al., 2023; Gloeckler et al., 2024).
@@ 3 Methods
3.1 Global solution from single-step transitions
Direct estimation of the global target distribution ($p({\bm{\theta}}|{\bm{x}}^{0:T})$ in NPE or $p({\bm{x}}^{0:T}|{\bm{\theta}})$ in NLE; Sec. 2.2) often requires a large number of simulations of entire time series, leading to high computational costs. We mitigate the problem of expensive simulation calls by leveraging the Markov factorization in Equation 1.
Since the transition probability $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ encapsulates all relevant information about the parameters, we expect that, given sufficiently many single-step transitions, it should be possible to infer the global target from observations of arbitrary time lengths. Building on this insight, we propose a general framework that focuses on solving the local inverse problem. In this framework, we aim to estimate a local target $s_{\text{local}}$ using single-step transition simulation data $\mathcal{D}=\{(\theta_{i},\mathbf{x}_{i}^{t:t+1})\}_{i=1}$ (Alg. 1). Afterward, we estimate a global target by aggregating these local solutions (Alg. 2), using a composition rule (compose, Sec. 3.2.2). In the following, we apply the framework to NSE, NLE, and NRE.
@@ 3.2.1 Local score estimation
Here, $\tilde{p}({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$ is a posterior with ${\bm{x}}^{t}$ following any proposal distribution ${\bm{x}}^{t}\sim\tilde{p}({\bm{x}}^{t})$ and ${\bm{x}}^{t+1}$ the state transition ${\bm{x}}^{t+1}\sim p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ respectively (Appendix Sec. A.1). The factorization (3) implies that global posterior is fully characterized by $s({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})=\nabla_{\theta}\log\tilde{p}({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$. We can aim to estimate this quantity using only single-state transitions by minimizing the following loss
$\mathcal{L}(\phi)=\mathbb{E}_{a,{\bm{\theta}},{\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1}}\left[\lambda(a)||s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1})-\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{\theta}})||_{2}^{2}\right]$ (4)
given any proposal distribution ${\bm{x}}^{t}\sim\tilde{p}({\bm{x}}^{t})$. Consequently, we can learn a local score estimator, as outlined in Alg. 1, by empirically minimizing this loss given a dataset of single-step simulations. This result generalizes to higher-order Markov chains, which require proposing more than one past observation (Appendix A.2.3 for details).
There is lots of flexibility in the choice of the proposal distribution $\tilde{p}({\bm{x}}^{t})$ in theory. The only requirement is that the proposed state must be independent of the parameters involved in the current state transition being emulated (see Alg. 1 line 4, Sec. A.1). However, this choice can have significant practical implications given a finite simulation budget, as further investigated in Sec. 4. In particular, this choice makes it possible to allocate simulation resources to inference questions of interest. For instance, if the goal is to perform inference on a predefined set of observations, the proposal can be tailored to predominantly sample values from this set. In general, however, designing a good proposal can be challenging.
Geffner et al. (2023) also proposed a score factorization in the i.i.d. setting to handle data-sets with multiple independent observations. The factorization (3) can be regarded as an extension of this factorization from the i.i.d. case to the Markov setting. Similarly, we can extend our results to partially factorized approaches by amortizing over multiple transitions instead of just a single one (see Appendix Sec. A.2.4).

### v2

@@ abstract
Amortized simulation-based inference (SBI) methods train neural networks on simulated data to perform Bayesian inference. While this strategy avoids the need for tractable likelihoods, it often requires a large number of simulations and has been challenging to scale to time series data. Scientific simulators frequently emulate real-world dynamics through thousands of single-state transitions over time. We propose an SBI approach that can exploit such Markovian simulators by locally identifying parameters consistent with individual state transitions. We then compose these local results to obtain a posterior over parameters that align with the entire time series observation. We focus on applying this approach to neural posterior score estimation but also show how it can be applied, e.g., to neural likelihood (ratio) estimation. We demonstrate that our approach is more simulation-efficient than directly estimating the global posterior on several synthetic benchmark tasks and simulators used in ecology and epidemiology. Finally, we validate scalability and simulation efficiency of our approach by applying it to a high-dimensional Kolmogorov flow simulator with around one million data dimensions.
@@ 1 Introduction
Numerical simulations are a central approach for tackling problems in a wide range of scientific and engineering disciplines, including physics (Brehmer & Cranmer, 2022; Dax et al., 2021), molecular dynamics (Hollingsworth & Dror, 2018), neuroscience (Gonçalves et al., 2020) and climate science (Watson-Parris et al., 2021). Simulators often include at least some parameters that cannot be measured experimentally. Inferring such parameters from observed data is a fundamental challenge. Bayesian inference provides a principled approach to identifying parameters that align with empirical observations (Gelman et al., 2020). Standard algorithms for Bayesian inference, such as Markov Chain Monte Carlo (MCMC) (Gilks et al., 1995) and variational inference (Beal, 2003), generally require access to the likelihoods $p({\bm{x}}|{\bm{\theta}})$. However, for many simulators, directly evaluating the likelihood remains intractable, rendering conventional Bayesian approaches inapplicable. Yet, generating synthetic data ${\bm{x}}\sim p({\bm{x}}|{\bm{\theta}})$ is feasible for numerical simulators. Simulation-based inference (SBI) methods offer a powerful alternative to perform Bayesian inference for such simulator models with intractable likelihoods (Cranmer et al., 2020).
@@
Fig 1: Illustration of Factorized Neural Score Estimation (FNSE). The goal is to perform parameter inference on a full time series model. The training process uses a smaller subsets of single-state transitions initialized at arbitrary proposal $\tilde{p}({\bm{x}}^{t})$, with parameters sampled from a prior distribution. During inference, the time series is divided into single state transitions, and each state transition is evaluated by the neural network to estimate local posterior scores. These local estimates are then aggregated to form a global approximation, which is subsequently used to sample from the overall posterior distribution. Here, $a$ denotes the diffusion time, and ${\bm{\theta}}_{a}$ is the associated noisy parameter.
@@
Classical SBI methods, like Approximate Bayesian Computation (ABC) (Beaumont et al., 2002) and synthetic likelihoods (Wood, 2010), struggle to effectively scale to high-dimensional simulations. To address this, SBI methods have been developed, which train neural networks to represent likelihoods (Papamakarios et al., 2019; Glöckler et al., 2022; Boelts et al., 2022), likelihood ratios (Durkan et al., 2020; Hermans et al., 2020; 2022; Miller et al., 2022), posteriors (Papamakarios & Murray, 2016; Lueckmann et al., 2017; Greenberg et al., 2019; Deistler et al., 2022; Geffner et al., 2023; Wildberger et al., 2023; Sharrock et al., 2024) or target several properties at once (Radev et al., 2023; Gloeckler et al., 2024b). These methods allow for parameter inference without requiring additional simulations after training, making them more efficient than traditional approaches (Lueckmann et al., 2021): they effectively amortize the cost of the simulation, and/or the full inference approach (Gershman & Goodman, 2014; Le et al., 2017).
@@
However, applying these neural amortized methods to time series simulations can be challenging due to the high computational cost of repeated simulator calls. Running numerous time series simulations—possibly with varying sequence lengths—can be computationally prohibitive or at least wasteful. Instead, it seems advantageous to exploit the temporal structure of these simulators: Many—if not most—scientific simulators for time series data are based on (stochastic) differential equations that model dynamics of a system through state transitions over time (e.g., Sedlmeier et al. (2016); Goswami et al. (2006); Strauss & Effenberger (2017); Sirur et al. (2016)), or directly model processes which are iteratively updated at each time-step. As a consequence, such simulators have an inherently Markovian structure, which can be leveraged for efficient inference! In particular, we can approach the problem on the level of single-state transitions, a simpler task that can be tackled with fewer simulations. For instance, in the Kolmogorov Flow simulator (Sec. 4.4), the simulation outputs reach nearly one million dimensions for a time series of length hundred, which is out of range for most direct SBI approaches. In contrast, the dimensionality of single state transitions is only a few thousand.
@@
In this work, we propose a method for efficient simulation-based inference for Markovian time series simulators. Unlike other neural SBI methods that require long sequences for training, our approach locally estimates parameters consistent with a single state transition. By aggregating these local solutions, we recover a global posterior approximation for time series of arbitrary length, significantly reducing simulation costs (Fig. 1). We focus on applying our strategy to posterior score-based approaches (FNSE, Factorized Neural Score Estimation) and additionally apply it to likelihood(-ratio) estimation methods (FNLE and FNRE). A related challenge has been addressed in prior studies  (Geffner et al., 2023; Linhart et al., 2024; Boelts et al., 2022) which had focused on efficient inference of i.i.d. (independently and identically distributed) observations. We here show that these approaches can be extended to Markovian simulators—arguably the dominant model-class for mechanistic simulators for time-series—and empirically evaluate our approach on a series of synthetic benchmark tasks, as well as simulators from ecology and epidemiology. Finally, we demonstrate the scalability and efficiency of this approach on the high-dimensional Kolmogorov flow problem.
@@ 2.1 Problem setting
To simplify the exposition, we constrain the initial condition to be independent of the parameter, $p({\bm{x}}^{0}|{\bm{\theta}})=p({\bm{x}}^{0})$, and assume that the simulation transition $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ does not depend on time $t$. We relax both of these assumptions in Appendix A.2, and also address the general case of higher-order Markov chains in Appendix A.2.3. We assume one can sample from the state-transition function ${\bm{x}}^{t+1}\sim\mathcal{T}({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$, but not evaluate its probability density.
@@ 2.2 Amortized methods for simulation-based inference
Neural Posterior Estimation (NPE): Neural Posterior Estimation (NPE; Papamakarios & Murray (2016); Lueckmann et al. (2017); Greenberg et al. (2019); Wildberger et al. (2023)) methods typically use a conditional neural density estimator, e.g., a normalizing flow (Rezende & Mohamed, 2015; Papamakarios et al., 2017; Durkan et al., 2019) to approximate the target posterior distribution. The model $q_{\phi}({\bm{\theta}}|{\bm{x}}^{0:T})$ is trained via maximum likelihood to estimate $p({\bm{\theta}}|{\bm{x}}^{0:T})$ from a dataset of parameter-data pairs $\{({\bm{\theta}}_{i},{\bm{x}}_{i}^{0:T})\}_{i=1}$, where $({\bm{\theta}}_{i},{\bm{x}}_{i}^{0:T})\sim p({\bm{\theta}})p({\bm{x}}^{0:T}|{\bm{\theta}})$. For observations of inhomogeneous size, NPE often incorporates an embedding network. Embedding networks can be designed as recurrent networks (Lueckmann et al., 2017), permutation-invariant architectures (Chan et al., 2018; Radev et al., 2020), or attention-based models (Schmitt et al., 2023) to respect known invariances in the data. At inference time, NPE yields an amortized approximation of the posterior for any observation ${\bm{x}}_{o}^{0:T}$ therefore offering rapid evaluations or samples of the posterior.
Neural likelihood(-ratio) estimation (NLE, NRE): Neural Likelihood(-Ratio) Estimation (NLE (Papamakarios et al., 2019), NRE (Hermans et al., 2020; 2022; Miller et al., 2022)) methods do not directly approximate the posterior distribution but instead train a surrogate $q_{\psi}({\bm{x}}_{i}^{0:T}|{\bm{\theta}})$ for the likelihood $p({\bm{x}}_{i}^{0:T}|{\bm{\theta}})$ or likelihood ratio $p({\bm{x}}_{i}^{0:T}|{\bm{\theta}})/p({\bm{x}}_{i}^{0:T})$. Once the likelihood is accessible via the surrogate, standard techniques such as MCMC or variational inference can be applied to perform inference. These approaches circumvent the problem of not having access to the likelihood, but still leaves us with the drawbacks of standard inference methods, i.e., slow sampling and potential failure modes such as robustly handling multimodality (Glöckler et al., 2022). Nonetheless, they also inherit the flexibility of standard inference techniques, such as the ability to handle multiple independent (or conditionally independent) observations by simply multiplying the likelihood terms (Boelts et al., 2022).
@@
This recently proposed approach has been highly successful across various tasks (Geffner et al., 2023; Sharrock et al., 2024; Gloeckler et al., 2024b). It offers a trade-off between the efficiency of the static NPE method and the flexibility of slower but more flexible NL(R)E method. By enabling feasible post-hoc modifications and composability through appropriate adjustments to the backward diffusion process, it bridges the gap between these two approaches (Geffner et al., 2023; Gloeckler et al., 2024b).
@@ 3 Methods
3.1 Global inference from single-step transitions
Direct estimation of the global target distribution (i.e., the posterior $p({\bm{\theta}}|{\bm{x}}^{0:T})$ in NPE or the likelihood $p({\bm{x}}^{0:T}|{\bm{\theta}})$ in NLE; Sec. 2.2) often requires a large number of simulations of the entire time series, leading to intractable, or at least very expensive, estimation problems. We mitigate the problem of expensive simulation calls by leveraging the Markov factorization of the forward model: Since the simulator is completely specified by its state-transition probabilities, $p({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$, these also contain all relevant information about the parameters. It is therefore possible to recover the global target distribution (e.g., the posterior) for a time series of variable size given a dataset of single-step transitions. Building on this observation, we aim to estimate local target scores $s_{\text{local}}$ using single-step transition simulation data $\mathcal{D}=\{(\theta_{i},\mathbf{x}_{i}^{t:t+1})\}_{i=1}^{N}$ (Alg. 1, Sec. 3.2.1). Afterwards, we recover the global target distribution by aggregating these local solutions using a composition rule (compose, Alg. 2, Sec. 3.2.2). In the following, we apply this approach to NSE, NLE, and NRE (Sec. 3.23.3) and discuss the design of the required proposal $\tilde{p}({\bm{x}}^{t})$ (Sec. 3.4) for sampling single-state transitions.
@@ 3.2.1 Local score estimation
Here, $\tilde{p}({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$ is a posterior with ${\bm{x}}^{t}$ following any proposal distribution ${\bm{x}}^{t}\sim\tilde{p}({\bm{x}}^{t})$ and ${\bm{x}}^{t+1}$ following the state transition ${\bm{x}}^{t+1}\sim\mathcal{T}({\bm{x}}^{t+1}|{\bm{x}}^{t},{\bm{\theta}})$ (Appendix Sec. A.1). The factorization (3) implies that the global posterior is fully characterized by $s({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})=\nabla_{\theta}\log\tilde{p}({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$. We can estimate this quantity using only single-state transitions by minimizing the loss
$\mathcal{L}(\phi)=\mathbb{E}_{a,{\bm{\theta}},{\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1}}\left[\lambda(a)||s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1})-\nabla_{{\bm{\theta}}_{a}}\log p({\bm{\theta}}_{a}|{\bm{\theta}})||_{2}^{2}\right],$ (4)
given a proposal distribution ${\bm{x}}^{t}\sim\tilde{p}({\bm{x}}^{t})$. As a result, we can learn a local score estimator by empirically minimizing this loss given a dataset of single-step simulations (Alg. 1). For a globally amortized posterior approximation, the proposal distribution must at least satisfy two properties: (i) it must have support at any point ${\bm{x}}^{t}$ the simulator can attain, (ii) the proposed point ${\bm{x}}^{t}$ must be independent of the parameter ${\bm{\theta}}$ involved in the transition to ${\bm{x}}^{t+1}$ to ensure Eq. 3 remains valid (Appendix Sec. A.1). These constraints leave quite some flexibility in the choice of the proposal. Its choice will have a significant impact, given a finite simulation budget (Sec. 3.4). Our approach can readily be extended to higher-order Markov chains, which require proposing multiple past observations (Appendix Sec. A.2.3). Between the two extreme cases of learning from single-state transitions, or full time series, there is a continuum of approaches, similarly to what was shown in Geffner et al. (2023) for independent observations. We can extend our results to amortizing over any finite number of transitions, leading to a partially factorized approach (Appendix Sec. A.2.4).